Solve the equation using the quadratic formula.
2x^2 = 30

Answer by dabanfield(803)

Solve the equation using the quadratic formula.
2x^2 = 30

Put the equation in standard form:
a*x^2 + b*x + c = 0
So we have:
2*x^2 + 0*x - 30 = 0
So a=2, b=0 and c = -30 in the quadratic formula.
Solve for x.